If n < 30 and σ İs unknown, then the 100(1-α)% confidence interval for a population mean is_ 7. assuming the population is normally distributed xtra 21 1(degrees of freedom = n-2) a. 1 (degrees of freedom-n-1) c. t(degrees of freedom-n-1) 1 d. x ± te |--| (degrees of freedom n-2) 1 (degrees of freedom = n-1) If n interval when the level of confidence is 99% 8. 16, then find the number ta2 needed in the construction of a confidence a. 2.583 b. 2.602 c. 2.921 d. 2.947 e. 2.977
Answer #1

7. when n < 30 and sigma is unknown we perform a t-test.

100(1-alpha)% confidence interval for the population mean:

Ans: b. T t to/2(degree of freedom n-1)

8. n = 16

df = n-1 = 15

100-99 = 1% 0.01

Using excel function : = T.INV.2T(0.01,15)

a/2-2.947

Answer d.